A line which makes equal angle on the lines x+y=1 and y=x,then slope of the line is/are

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
indirectly we have to find angle bisector of given lines
 
x+ y – 1 / sqrt(2) = ± ( x – y/ sqrt(2))
 
hence first line is
 
x+ y – 1 = x – y
2y = 1
 
second line is
x + y - 1 = y – x
 
2x = 1